" what are the dimensions and centimeters?"
The length of a rectangle is 2 centimeters less than its width.What are the dimensions of the rectangle if its area is 168 square centimeters?
The length of the rectangle is____ centimeters
The width of the rectangle is ____ centimeters

Answer by Jresend18(1)

x^2-2x=168
x^2-2x-168=0
(x+12)(x-14)=0
x=-12,x=14
length=-12
width=14
solutions: {-12,14}